OpenStudy (anonymous):

In one week, Kean ate a total of 8 ounces of grains at breakfast, fourteen ounces of grains at lunch, and sixteen ounces of grains at dinner. Which of the following shows a simplified ratio of his servings of grain at breakfast to his servings of grain at lunch?

OpenStudy (anonymous):

the choices says A. 4:7 B.4/19 C.7TO 4 D.4:15

OpenStudy (anonymous):

Hold on.

OpenStudy (anonymous):

K.

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

can u show me how u got that please?

OpenStudy (anonymous):

8 + 14 + 16 = 38 then; divide the breakfast servings to the total servings

OpenStudy (anonymous):

then simplify 8/38=4/19

OpenStudy (anonymous):

understand it now @nataliedhatshortiee ?

OpenStudy (anonymous):

yes thanku ur aswsome :)

OpenStudy (anonymous):

If you need more help just tag me and ill answer/help you